OpenStudy (yanasidlinskiy):

Help..... Simplify the expression. 12P5

hartnn (hartnn):

\(\Large ^nP_r = \dfrac{n!}{(n-r)!}\)

hartnn (hartnn):

here, n = 12, r = 5 plug in values!

OpenStudy (yanasidlinskiy):

\[=\frac{ 12! }{ (12-5)! } = \frac{ 12 }{ (7)! }\] What would I do after that?

hartnn (hartnn):

you know how n! gets calculated ?

OpenStudy (yanasidlinskiy):

umm...idk..

hartnn (hartnn):

\(\Large n! = n \times (n-1)\times (n-2)\times ... 3 \times 2\times 1 \\ eg: 4! = 4\times 3\times 2\times 1\)

OpenStudy (anonymous):

7!=7*6*5*4*3*2*1

hartnn (hartnn):

another example : \(7! = 7 \times 6\times 5\times 4\times 3\times 2\times 1\)

hartnn (hartnn):

oh and the numerator is 12! isn't it ?

hartnn (hartnn):

\(=\frac{ 12! }{ (12-5)! } = \frac{ 12! }{ (7)! }\)

OpenStudy (yanasidlinskiy):

Yes. It is. I am going to solve this and you tell me if I'm right. Ok?

hartnn (hartnn):

sure

OpenStudy (yanasidlinskiy):

\[\frac{ 12*11*10*9*8*7*6*5*4*3*2*1 }{ 7*6*5*4*3*2*1 }=95,040\] Am i right?

hartnn (hartnn):

thats absolutely correct! :)

OpenStudy (yanasidlinskiy):

Yay!!:) Wow...So proud....When I submit...my exam...I will tell you what score i got!!:) lol!!

hartnn (hartnn):

here's a neat trick to do these faster, say we have 6! / 3! we know that 3! = 3.2.1 and 6! = 6.5.4.3.2.1 see anything common ? yes! 3.2.1 so, 6! = 6.5.4.3! so, 6!/3! is just 6.5.4 :)

hartnn (hartnn):

similarly, your 12!/ 7! will simply be 12.11.10.9.8

OpenStudy (yanasidlinskiy):

Wait.....I have one more question..... A number cube is rolled with these results: 64 ones, 67 twos, 73 threes, 59 fours, 72 fives, and 71 sixes. What is the experimental probability of rolling an even number? Write your answer as a percent, to the nearest tenth of a percent. A. 51.9% B. 48.5% C. 53.6% D. 46.8%

OpenStudy (yanasidlinskiy):

Nice!:) Helps e a lot....on my second question i think it would be either B or D.

hartnn (hartnn):

Experimental Probability = favorable events/ total events = (# of rolls producing even numbers)/ (Total # of rolls) Total # of rolls = 64+67+...+71 =... for # of rolls producing even numbers, you consider , 2,4,6 only = 67+59+71 = ... then just divide :)

OpenStudy (yanasidlinskiy):

So i would add them before i divide...right?

hartnn (hartnn):

yesss

hartnn (hartnn):

see if you get 0.485 , which is 48/5% :)

hartnn (hartnn):

**48.5 %

OpenStudy (yanasidlinskiy):

Um...ok. no. i got like 33.833..I added 67, 64...etc and then added 2,4,6...seperately..then i divided 406 by 12..Thanks for your help though..

OpenStudy (yanasidlinskiy):

I got a 100% on my probability test!!!:)